calculus 3 please fast

AT233 t1 Summer 2018 6. Determine if the following lines are parallel, intersecting or skew
Answer #1

Solution:

L2 are (.-2. 3〉&く10.4.6) Th e direction vectors parallel to Li an

The lines are not parallel because their vectors are not proportional

If the lines were to intersect we would have three equations in two unknowns s and t

4 + 5-10s

-2t=6+4s

1+3t=4+6s

Using the first two equations to solve for s and t we get

0 19 s=__ and t =-

Using the second two equations to solve for s and t we get

s-1 and t-1

Thus the lines do not intersect because different values of t and s are obtained

for solving the first and second, and then the second and third equations.

Since the lines are not parallel and they do not intersect, then they must be skew